Liam:

Which of the following is not a direction vector for the plane 3x-2y-z+4=0? a) (3,-2,-1) b) (0,-2,4) c) (2,2,2) (2,4,-2) How do you know when its not a direction vector?

Liam:

@Vocaloid

Vocaloid:

disclaimer: had to look this one up but I believe you would plug in the x,y, and z value for each coordinate the +4 can be neglected [not entirely sure why, I think it has something to do with the fact that we are only considering direction] 3x-2y-z = 0 then just see which answer choice makes this statement true

Vocaloid:

idk when they'll be on but @hero @Angle

Vocaloid:

oh wait, it says which one is ~not~ a direction vector so you would pick the vector that makes the statement false

Liam:

Yeah, so I got this: (3,-2,-1) So option A is off.

Liam:

So let me try plugging in now

Liam:

So hold up. B and C make the statement false. :|

Vocaloid:

hm? 3x-2y-z = 0 B gives (0,-2,4) --> 3(0) - 2(-2) - 4 = 0 C gives (2,2,2) --> 3(2) - 2(2) - 2 = 0 so I believe C and B are both still direction vectors going back to choice A I end up with 3(3) - 2(-2) - 1(-1) which is 9 + 4 + 1 = 14 so I'm actually inclined to go w/ choice a

Liam:

Okay, but if we do option d the statement is also not correct. It gives us 4=0. So?

Liam:

And I'm pretty sure option A can't be correct.

Vocaloid:

hm. I can ask somebody else's opinion b/c I'm still not entirely sure how to do this

Liam:

Okay, please do. Tysm!

Vocaloid:

@liam alright I just got back for a direction vector, you consider every plane that is parallel to the original plane, not just the original plane, so the +4, for the sake of calculations, can be disregarded so instead of calculating 3x-2y-z+4=0 you can simply use 3x - 2y - z = 0 and plug in the coordinates of each vector